# Affordance-based Robot Manipulation with Flow Matching

### v3

@@ Abstract
We present a framework for assistive robot manipulation, which focuses on two fundamental challenges: first, efficiently adapting large-scale models to downstream scene affordance understanding tasks, especially in daily living scenarios where gathering multi-task data involving humans requires strenuous effort; second, effectively learning robot trajectories by grounding the visual affordance model. We tackle the first challenge by employing a parameter-efficient prompt tuning method that prepends learnable text prompts to the frozen vision model to predict manipulation affordances in multi-task scenarios. Then we propose to learn robot trajectories guided by affordances in a supervised Flow Matching method. Flow matching represents a robot visuomotor policy as a conditional process of flowing random waypoints to desired robot trajectories. Finally, we introduce a real-world dataset with 10 tasks across Activities of Daily Living to test our framework. Our extensive evaluation highlights that the proposed prompt tuning method for learning manipulation affordance with language prompter achieves competitive performance and even outperforms other finetuning protocols across data scales, while satisfying parameter efficiency. Learning multi-task robot trajectories with flow matching policy also leads to consistently better results than alternative behavior cloning methods, including marginally better generalization performance and prominently faster inference than diffusion policy with DDPM. Our framework seamlessly unifies affordance model learning and trajectory generation with flow matching for robot manipulation. https://hri-eu.github.io/flow-matching-policy/.
1 Introduction
@@
Recent advances in vision-language models (VLMs) present unprecedented opportunities to solve robot manipulation problems. Attempts in the field have focused on three primary aspects: 1) End-to-end learning manipulation from scratch. These approaches (Padalkar et al., 2023) make the least assumptions on tasks and are formulated in language-image-to-action prediction models. 2) Off-the-shelf-vision-language models for robot manipulation. Such line of works have explored using pre-trained VLMs with prompt engineering in various contexts of robot motion learning, including reward design for reinforcement learning (Ma et al., 2023), python coding (Liang et al., 2023), joint actions (Wang et al., 2023), etc. 3) Intermediate substrate to bridge high-level language-image instructions and low-level robot policies. These works usually introduce some form of prior derived from human knowledge as an intermediate stage, including affordances (Huang et al., 2023), pre-trained visual representations (Xiao et al., 2022), primitive skills (Ingelhag et al., 2024), etc. In this paper, we follow the third line of work to unify an affordance model and low-level robot policies. which helps to alleviate the sample inefficiency problem of end-to-end learning from scratch.
@@
Figure 1: The proposed framework of unifying affordance model learning and trajectory generation for robot manipulation. Given the same visual scene but different language instructions, the model first extracts instruction-relevant manipulation affordances. This is achieved through a prompt tuning method that prepends learnable text-conditioned prompts in a frozen vision foundation model. Then a flow matching policy is proposed to transform the random waypoints to the desired trajectories, guided by task-relevant affordances.
@@
Extracting affordance knowledge has long inspired the robot community (Xiao et al., 2022). Humans heavily rely on visual affordances to efficiently perform day-to-day tasks across environments. The concept of affordance has been introduced in (Gibson, 2014), referring to the ability to perform certain actions with objects in the context of a given scene. In this work, affordance focuses on the areas of an object that permit or enable specific actions. Affordance models allow more knowledge transfer across different objects with similar affordances, and more intuitive interaction with environments, compared to other intermediate substrates, including keypoints or rewards. We particularly focus on affordance learning in multi-task scenarios with text prompting. As shown in Fig. 1, given the same visual scene but with different language instructions, we aim to extract different affordances for robot policy learning through our proposed model. To leverage the ability of pre-trained foundation models while avoiding expensive computational constraints, recent works have explored parameter-efficiently fine-tuning (PEFT) large vision-language models for various visual recognition applications (Jia et al., 2022). PEFT could be mainly categorized into two main groups: adapter-based methods (e. g., LoRA (Hu et al., 2021)) and prompt-based methods (Liu et al., 2023). One representative line of such research has concentrated on prompt tuning methods, which prepend learnable prompts to the input of a large frozen pre-trained model and optimize them via gradients during finetuning. Studies on randomly-generalized trainable prompts (Jia et al., 2022) for universal use or condition-admitted prompt variables (Sohn et al., 2023) for better specific task performance have been both explored. It has shown that prompt tuning could match the performance of full finetuning but with substantially less parameter storage in various domains, including visual tracking (Zhu et al., 2023) and cross-domain tasks such as language-dance assessment (Zhong and Demiris, 2024). Inspired in part by the notion of human cognitive penetrability mechanism (Maier and Abdel Rahman, 2019) that uses linguistic knowledge to tune ongoing visual processing, we target to incorporate learnable text-conditioned prompts into any vision foundation model while keeping it frozen, preserving its visual understanding capabilities, to learn instruction-relevant manipulation affordance.
@@
The subsequent challenge involves deploying the visual affordance across various robot manipulation learning paradigms. From the traditional behavior cloning with convolutional networks (Zhang and Demiris, 2022) to transformer-based learning structures (Shridhar et al., 2023), extensive research has modeled robot action trajectories from visual scenes. A recent line of works builds on successes in diffusion models (Chi et al., 2023) to generate motion trajectories to capture multimodal action distributions. Flow Matching is another novel generative method. Sharing theoretical similarities with stochastic Denoising Diffusion Probabilistic Models (DDPM), flow matching aims to regress onto a deterministic vector field to flow samples toward the target distribution. It has proven that the simplicity of flow matching objectives allows favorable performance in stable training and generation quality compared to solving complex stochastic differential equations in DDPM. We extend flow matching to the robotics domain. As shown in Fig. 1 and Fig. 1, the proposed method would flow the random waypoints to the desired trajectories based on various affordances in a single flow matching policy.
Figure 2: Flow Matching policy representations. The flow, waypoint motion path, and vector field are presented to showcase the process of flow matching transforming random waypoints to the target trajectory from timestep 0 to 1.
@@
We also construct a real-world dataset with 10 tasks across Activities of Daily Living to test the proposed method. The novelty of our dataset is that it contains the same scenarios but with multi-task affordance and robot trajectories. Experimental evaluation on our dataset and other benchmarks empirically demonstrates that the prompt tuning method for learning affordances achieves performance competitive, and sometimes beyond other finetuning protocols across data scales, vision-language fusion architectures, and prompt variants. Furthermore, we showcase that flow matching policy attains favorable performance in stable training, generation quality and computational efficiency amongst competing methods of robot behavior cloning.
This is the first attempt to ground VLM-based affordance with flow matching for real-world robot manipulation. Note that our goal is not to achieve the state-of-the-art general robot manipulation performance, but instead to broadly explore a new paradigm of efficiently adapting VLMs for affordance learning, and robot policy for multimodal action distributions.
@@
    A novel formulation using Flow Matching for robot manipulation policy learning from visual affordances.
@@
    Empirical results show that flow matching policy leads to consistently better results than alternative behavior cloning methods, including marginally better generalization performance and prominently faster inference than diffusion policy with DDPM.
2 Related Work
2.1 Robot Learning from Demonstration
@@
Imitation learning has been a common paradigm for robots which requires simulated or real-world demonstration data collection (Zhang and Demiris, 2023). To improve data efficiency, extensive works have been proposed to learn robot policies on the top of visual representations (Liu et al., 2024) instead of end-to-end raw images (Goyal et al., 2023). Keypoints or affordance heatmaps are often used to provide contact information as visual representation (Liu et al., 2024). This paper concentrates on using affordances to guide the low-level robot manipulation. In terms of network architectures for robot learning, prior works have successfully investigated convolutional networks (Zhang and Demiris, 2022), Transformers (Shridhar et al., 2023), generative adversarial networks (Ho and Ermon, 2016), Energy-Based Models (Florence et al., 2022), etc. However, the collected data is usually expected to be non-convex and multi-modal due to the variability in human demonstrations. Recent works have addressed this problem by reformulating the robot policy as a generative process. Diffusion policy (Chi et al., 2023) has emerged as a powerful class of generative models for behavior cloning by representing a robot’s visuomotor policy as a conditional denoising diffusion process. In this work, we investigate Flow Matching (Lipman et al., 2022), a novel generative model that has demonstrated its superiority in image generation, but is much less explored in robotics domains.
2.2 Parameter-Efficient Finetuning
@@
Instruction-aware vision encoding (Gupta et al., 2022) has been extensively studied for various language-vision fusing tasks (Radford et al., 2021). Given the dominance of large-scale vision-language models, many approaches have been proposed to efficiently finetune a frozen pretrained model for different downstream tasks to speed up training and reduce memory. Two representative methods among them are adapters and prompting. The first type of research varies depending on the adapter that could add extra lightweight modules (Gao et al., 2024) or express the weight updates as a low-rank decomposition of the weight matrix (Hu et al., 2021). Another line of work focuses on prompting (Liu et al., 2023), which originally primes a frozen pretrained language model for downstream tasks by including a text prompt. Recent works on prompt tuning (Liu et al., 2021) treat prompts as continuous vectors and compute their gradients with backpropagation during training. The extension of prompt tuning to vision tasks has gained massive success. Visual prompt tuning (Jia et al., 2022) has manipulated visual prompts to steer models in arbitrary vision tasks. Other works (Sohn et al., 2023) have further explored extending the random learnable prompts to condition-based prompts (e. g., class, instance) that are hence less universal but more accurate. Inspired by its recent success, we extend the prompt tuning technologies to address the challenge of adapting large pretrained vision-language models to affordance learning for robot manipulation. The intuition is clear: if the model understands the posed text instruction and the inherent context, it should extract visual affordances that directly correspond to the relevant image aspects. Our method achieves the above goal by integrating learnable text-conditioned prompts into a large vision encoder, while keeping it frozen to preserve visual understanding capabilities.
2.3 Flow Matching in Robotics
Despite its recent progress in image generation (Albergo and Vanden-Eijnden, 2022), the application of flow matching in robotics domains remains underexplored. Few prior studies (Braun et al., 2024; Hu et al., 2024; Rouxel et al., 2024) have successfully modeled transformations that smoothly move random samples towards target robot actions. These works are nevertheless limited to simulation tasks or conditioned on state-based input. We propose the flow matching policy to learn multi-task robot behavior from raw images with visual affordances in a single supervised policy.
3 Methods
3.1 Prompt Tuning for Affordance Learning
@@
Given a pre-trained vision transformer, our objective is to learn a set of text-conditioned prompts to maximize the likelihood of correct affordance labels, as shown in Fig. 3. Only the prompt-related layers and the decoder are being updated during the training, while the vision transformer remains frozen. Inspired by Vision Prompt Tuning (Jia et al., 2022), we also propose shallow and deep network architectures.
@@ 3.1.2 Deep Architecture
3.1.3 Implementation Details
Our goal is to integrate textual representations into any vision encoder while keeping it frozen, preserving its visual understanding capabilities. Thus, we have chosen the most basic vision backbone, a pretrained ViT-B-16 transformer. The text layer adopts the classic CLIP setup, including Multiheaded Self-Attention, Feed-Forward Networks with LayerNorm and residual connections. As suggested by MAE (He et al., 2022), the decoder is only used for downstream tasks and could be flexible and lightweight. Thus we use one single transformer decoder layer.
We adopt the L2 Mean Squared loss between the predicted and ground truth affordances for network training. We add positional embeddings to all the image and language tokens to preserve the positional information. In the subsequent experiments, we will further study multiple model variants, including text and vision fusion, prompt depth, pretrained weights for vision transformer, etc.
3.2 Flow Matching Policy
We build the robot behavioral cloning policy as a generative process of Flow Matching, which constructs a flow vector that continuously transforms a source probability distribution toward a destination distribution. Flow Matching leverages an ordinary differential equation to deterministically mold data distribution, contrasting with diffusion policy which is based on a stochastic differential equation through introducing noise.
3.2.1 Flow Matching Model
@@
where $\bm{x}\sim p_{t}(\bm{x}|\bm{z})$, $t\sim\mathcal{U}[0,1]$ (uniform distribution). Flow matching aims to regress $\mathbf{u}_{t}(\bm{x}|\bm{z})$ with a time-dependent vector field of flow $\mathbf{v}_{t}(\bm{x},\bm{\theta})$ parameterized as a neural network with weights $\bm{\theta}$. $\mathbf{u}_{t}(\bm{x}|\bm{z})$ can be further simplified as:
@@
$p_{0}$ represents a simple base density at time $t=0$, $p_{1}$ denotes the target complicated distribution at time $t=1$, $\bm{x}_{0}$ and $\bm{x}_{1}$ are the corresponding samplings. $\mathbf{v}_{t}(\bm{x},\bm{\theta})$ could be described as
@@
We define $\bm{x}_{t}$ as the linear interpolation between $\bm{x}_{0}$ and $\bm{x}_{1}$ with respect to time $\bm{x}_{t}=t\bm{x}_{1}+(1-t)\bm{x}_{0}$, following the Optimal Transport theory (Peyré et al., 2019). And $v_{\bm{\theta}}$ is a network of the flow model. Thus Equation (1) could be reformatted as
@@
Figure 4: Framework of Flow Matching Policy. (a) At each time step, the policy predicts flow vectors for each waypoint conditioned on visual observation data as well as the waypoints at the current timestep. (b-c) Visualization of the inference phase using Flow matching to transform random waypoints (green) to the target action trajectory (purple) from timestep 0 to 1. The lines in (b) denote the flow vectors.
3.2.2 Flow Matching for Visuomotor Policy Learning

### v4

@@ Abstract
We present a framework for assistive robot manipulation, which focuses on two fundamental challenges: first, efficiently adapting large-scale models to downstream scene affordance understanding tasks, especially in daily living scenarios where gathering multi-task data involving humans requires strenuous effort; second, effectively learning robot action trajectories by grounding the visual affordance model. We tackle the first challenge by employing a parameter-efficient prompt tuning method that prepends learnable text prompts to the frozen vision model to predict manipulation affordances in multi-task scenarios. Then we propose to learn robot action trajectories guided by affordances in a supervised flow matching method. Flow matching represents a robot visuomotor policy as a conditional process of flowing random waypoints to desired robot action trajectories. Finally, we introduce a real-world dataset with 10 tasks across Activities of Daily Living to test our framework. Our extensive evaluation highlights that the proposed prompt tuning method for learning manipulation affordance achieves competitive performance and even outperforms some other finetuning protocols across data scales, while satisfying parameter efficiency. Learning multi-task robot action trajectories with flow matching leads to consistently favorable results in several robot manipulation benchmarks than some alternative behavior cloning methods. This includes more stable training and evaluation, and noticeably faster inference, while maintaining comparable generalization performance to diffusion policy, where flow matching performs marginally better in most cases. Our framework seamlessly unifies affordance learning and action generation with flow matching for robot manipulation. https://hri-eu.github.io/flow-matching-policy/
I Introduction
@@
Recent advances in vision-language models (VLMs) present unprecedented opportunities to solve robot manipulation problems. Attempts in the field have focused on three primary aspects: 1) End-to-end learning manipulation from scratch. These approaches [35] make the least assumptions on tasks and are formulated in language-image-to-action prediction models. 2) Off-the-shelf-vision-language models for robot manipulation. Such line of works have explored using pre-trained VLMs with prompt engineering in various contexts of robot motion learning, including reward design for reinforcement learning [31], python coding [26], joint actions [47], etc. 3) Intermediate substrate to bridge high-level language-image instructions and low-level robot policies. These works usually introduce some form of prior derived from human knowledge as an intermediate stage, including affordances [22], pre-trained visual representations [48], primitive skills [23], etc. In this paper, we follow the third line of work to unify an affordance model and low-level robot policies, which helps to alleviate the sample inefficiency problem of end-to-end learning.
@@
Figure 1: The proposed framework of unifying affordance map learning and action generation for robot manipulation. Given the same visual scene with different language instructions, the model first extracts instruction-relevant manipulation affordances. This is achieved through a prompt tuning method that prepends learnable text-conditioned prompts in a frozen vision foundation model. Then a flow matching policy is proposed to transform the random waypoints to the desired action trajectories, guided by task-relevant affordance maps.
Extracting affordance knowledge has long inspired the robot community [48]. Humans heavily rely on affordances to efficiently perform day-to-day tasks across environments. The concept of affordance has been introduced in [12], referring to the ability to perform certain actions with objects in the context of a given scene. In this work, we first focus on learning affordance maps related to the locations of object areas that permit or enable specific actions. Then, the knowledge of affordance maps is used to plan and execute appropriate physical interactions with objects with our robot learning policy.
@@
We particularly concentrate on multi-task scenarios with text prompting. As shown in Fig. 1, given the same visual scene but with different language instructions, we aim to extract different affordances for robot policy learning through our proposed model. To leverage the ability of pre-trained foundation models while avoiding expensive computational constraints, recent works have explored parameter-efficiently fine-tuning (PEFT) large vision-language models for visual recognition applications [24]. PEFT could be mainly categorized into two main groups: adapter-based methods (e. g., LoRA [20]) and prompt-based methods [29]. One representative line of such research has concentrated on prompt tuning methods, which prepend learnable prompts to the input of a large frozen pre-trained model and optimize them via gradients during finetuning. Studies on randomly-generalized trainable prompts [24] for universal use or condition-admitted prompt variables [43] for better specific task performance have been both explored. It has shown that prompt tuning could match the performance of full finetuning but with substantially less parameter storage in various domains, including visual tracking [53] and cross-domain tasks such as language-dance assessment [52]. Inspired in part by the notion of human cognitive penetrability mechanism [32] that uses linguistic knowledge to tune ongoing visual processing, we target to incorporate learnable text-conditioned prompts into any vision foundation model while keeping it frozen, preserving its visual understanding capabilities, to learn instruction-relevant manipulation affordance maps.
@@
The subsequent challenge involves deploying the visual affordance across robot manipulation learning paradigms. From the traditional behavior cloning with convolutional networks [49] to transformer-based learning structures [42], extensive research has modeled robot action trajectories from visual scenes. A recent line of work builds on successes in diffusion models [3] with Denoising Diffusion Probabilistic Models (DDPM [19]) and Denoising Diffusion Implicit Models (DDIM [44]) to generate motion trajectories to capture multimodal action distributions. Flow Matching is another novel generative method. Sharing theoretical similarities with diffusion process, flow matching aims to regress onto a deterministic vector field to flow samples toward the target distribution. [27] has proven that the simplicity of flow matching objectives allows favorable performance in stable training and generation quality compared to solving complex stochastic differential equations in stochastic denoising diffusion process. We extend flow matching to the robotics domain. As shown in Fig. 1, the proposed method would flow the random waypoints to the desired action trajectories based on multi-task affordances in a single flow matching policy.
@@
We also construct a Real-world Activities of Daily Living (ADLs) dataset with 10 tasks. The novelty of our dataset is that it contains the same scenarios but with multi-task affordance and robot trajectories. Experimental evaluation on our dataset empirically demonstrates that the prompt tuning method for learning affordances achieves performance competitive, and sometimes beyond other finetuning protocols across data scales, vision-language fusion architectures, and prompt variants. Furthermore, we showcase that across several benchmarks, flow matching attains favorable performance in training stability, generation quality, and computational efficiency amongst competing methods of behavior cloning.
This is the first attempt to ground VLM-based affordance with flow matching for real-world robot manipulation. We have also first systematically evaluated robot manipulation with flow matching on several benchmarks, including various input representations, robot control types, and manipulation tasks. Our code has been released https://github.com/HRI-EU/flow_matching.
@@
    A novel formulation using flow matching for closed-loop 6D robot manipulation learning from various inputs, including visual affordances.
@@
    Empirical and extensive results show that flow matching leads to consistently favorable results than some alternative behavior cloning methods in various manipulation tasks. This includes more stable training and evaluation, and noticeably faster inference than diffusion policy, while maintaining comparable generalization performance to DDIM and slightly better performance than DDPM.
Note that our goal is not to achieve the state-of-the-art general robot manipulation performance, but instead to broadly explore a new paradigm of efficiently adapting VLMs for affordance learning, and robot policy for multimodal action distributions.
II Related Work
II-A Robot Learning from Demonstration
@@
Imitation learning has been a common paradigm for robots which requires simulated or real-world demonstration data collection [50]. To improve data efficiency, extensive works have been proposed to learn robot policies on the top of visual representations [28], such as keypoints or affordance heatmaps [28], instead of end-to-end raw images [13]. This paper concentrates on using affordances to guide the low-level robot manipulation. In terms of network architectures for robot learning, prior works have successfully investigated convolutional networks [49], Transformers [45], generative adversarial networks [18], Energy-Based Models [7], etc. However, the collected data is usually expected to be non-convex and multi-modal due to the variability in human demonstrations. Recent works have addressed this problem by reformulating the robot policy as a generative process. Diffusion policy [3] has emerged as a powerful class of generative models for behavior cloning by representing a robot’s visuomotor policy as a conditional denoising diffusion process. In this work, we investigate flow matching [27], a novel generative model that has demonstrated its superiority in image generation, but is much less explored in robotics domains.
II-B Parameter-Efficient Finetuning
@@
Instruction-aware vision encoding [15] has been extensively studied for language-vision fusing tasks [39]. Given the dominance of large-scale vision-language models, many approaches have been proposed to efficiently finetune a frozen pretrained model for downstream tasks to speed up training and reduce memory. Two representative methods among them are adapters and prompting. The first line of research varies depending on the adapter that could add extra lightweight modules [10, 20]. Other work focuses on prompting [29], which originally primes a frozen pretrained language model for downstream tasks by including a hard text prompt. Recent works on prompt tuning [30] treat soft prompts as continuous vectors and compute their gradients with backpropagation during training. The extension of prompt tuning to vision tasks has gained massive success. Visual prompt tuning [24] has manipulated visual prompts to steer models in arbitrary vision tasks.  [43] have explored extending the random learnable prompts to condition-based prompts that are less universal but more accurate. Inspired by its recent success, we extend the prompt tuning technologies to address the challenge of adapting large pretrained vision-language models to affordance learning for robot manipulation. The intuition is clear: if the model understands the posed text instruction and the inherent context, it should extract visual affordances that directly correspond to the relevant image aspects. Our method achieves the above goal by integrating learnable text-conditioned prompts into a large vision encoder, while keeping it frozen to preserve visual understanding capabilities.
II-C Flow Matching in Robotics
Despite its recent progress in image generation [1], the application of flow matching in robotics domains remains underexplored [21, 41]. Few prior studies have concentrated exclusively on certain robot scenarios for deploying flow matching, for pointcloud environment [4], Riemannian manifold [5], etc. We propose to use flow matching to learn multi-task robot behaviors from raw observations, including visual affordances obtained from a vision-language model, in a single supervised policy. We have first systematically evaluated robot manipulation with flow matching on several benchmarks, including various input representations, robot control types, and manipulation tasks.
III Methods
III-A Prompt Tuning for Affordance Map Learning
@@
Providing any type of pre-trained vision transformer, our objective is to learn a set of text-conditioned prompts to maximize the likelihood of correct affordance labels, as shown in Fig. 2. Only the prompt-related layers and the decoder are being updated during the training, while the vision transformer remains frozen. Inspired by Vision Prompt Tuning [24], we propose shallow and deep network architectures.
@@ III-A2 Deep Architecture
Figure 3: Framework of flow matching policy. (a) General formulation. At each time step, flow matching takes visual observation $\bm{o}$ (e. g., state-based inputs, RGB-D images, visual affordances) as input, and outputs robot actions (e. g., 6D robot end-effector trajectories, robot joint actions, gripper actions). (b-d) Visualization of the inference process of transforming random waypoints to target actions over time from 0 (green) to 1 (purple). Green lines in (b) denote the flow paths. Vector fields are shown in (c).
III-A3 Implementation Details
Our goal is to integrate textual representations into any vision encoder while keeping it frozen, preserving its visual understanding capabilities. Thus, we have chosen the most basic vision backbone, a pretrained ViT-B-16 transformer. The text layers adopt the classic CLIP setup [39], including Multiheaded Self-Attention, Feed-Forward Networks with LayerNorm and residual connections, and output 76 tokens. As suggested by MAE [17], the decoder is only used for downstream tasks and could be flexible and lightweight. Thus we use one single transformer decoder layer.
We adopt the L2 Mean Squared loss between the predicted and ground truth affordances for network training. The training parameters for the prompt tuning network include an image size of 224 × 224, AdamW with a learning rate of 1.5e-4 including Warmup with step-decay, and batch size of 256. We add positional embeddings to all the image and language tokens to preserve the positional information. In the subsequent experiments, we will further ablate multiple model variants, including text and vision fusion structures, prompt depth, pretrained weights for vision transformers, etc.
III-B Flow Matching Policy
We build the robot behavioral cloning policy as a generative process of flow matching, which constructs a flow vector that continuously transforms a source probability distribution toward a destination distribution. Flow matching leverages an ordinary differential equation to deterministically mold data distribution, contrasting with Denoising Diffusion Probabilistic Models (DDPM) that is based on a stochastic differential equation through introducing noise.
III-B1 Flow Matching Model
@@
where $\bm{x}\sim p_{t}(\bm{x}|\bm{z})$, $t\sim\mathcal{U}[0,1]$. Flow matching aims to regress $\mathbf{u}_{t}(\bm{x}|\bm{z})$ with a time-dependent vector field of flow $\mathbf{v}_{t}(\bm{x},\bm{\theta})$ parameterized as a neural network with weights $\bm{\theta}$. $\mathbf{u}_{t}(\bm{x}|\bm{z})$ can be further simplified as:
@@
$p_{0}$ represents a simple base density at time $t=0$, $p_{1}$ denotes the target complicated distribution at time $t=1$, $\bm{x}_{0}$ and $\bm{x}_{1}$ are the corresponding samplings. $\mathbf{v}_{t}(\bm{x},\bm{\theta})$ could be described as:
@@
where we define $\bm{x}_{t}$ as the linear interpolation between $\bm{x}_{0}$ and $\bm{x}_{1}$ with respect to time $\bm{x}_{t}=t\bm{x}_{1}+(1-t)\bm{x}_{0}$, following the Optimal Transport theory [37]. And $v_{\bm{\theta}}$ is a network of the flow model. Thus Equation (1) could be reformatted as
@@
III-B2 Flow Matching for Visuomotor Policy Learning